# Bounding the Treewidth of Outer $k$-Planar Graphs via Triangulations

### v1

@@ Abstract
treewidth, outer $k$-planar graphs, outer min-$k$-planar graphs, cop number, separation number
@@ 1 Introduction
Treewidth is an important structural parameter because the running time of many graph algorithms and algorithms for drawing graphs depend on the treewidth. Recently treewidth-based techniques have been successfully applied for convex drawings, that is, straight-line drawings where vertices lie on a circle. For example, Bannister and Eppstein [8] showed that if a graph admits a convex drawing with at most $k$ crossings in total, then its treewidth is bounded by a function of $k$. Furthermore, they used this fact to give, for a fixed $k$, a linear-time algorithm (via extended monadic second-order logic and Courcelle’s theorem [13]) to decide whether a convex drawing with at most $k$ crossings exists for a given graph. Chaplick et al. [12] generalized this to bundled crossings (a crossing of two bundles of edges such that in each bundle the edges travel in parallel, which is counted as one crossing). Another prominent class of graphs with convex drawings is the class of outer $k$-planar graphs, that is, graphs that admit a convex drawing where each edge is crossed at most $k$ times. Unlike $k$-planar graphs (without the restriction on the placement of vertices), the treewidth of outer $k$-planar graphs can be bounded by a function of $k$ only (see discussion below). Similarly to Bannister and Eppstein [8], Chaplick et al. [10] used this to test in linear time, for any fixed $k$, whether a given graph is full outer $k$-planar, i.e., whether it admits an outer $k$-planar drawing where no crossing appears on the boundary of the outer face.
@@
For constant $k$, better treewidth bounds are known. A folklore result is that outerplanar graphs ($k=0$) have treewidth at most $2$. Auer et al. [5, 6] showed that maximal outer $1$-planar graphs are chordal graphs and, therefore, have treewidth at most $3$, which is tight.
@@
A structurally similar type of result is known for $2$-layer $k$-planar graphs. A $2$-layer $k$-planar graph is a bipartite graph that admits a $k$-planar straight-line drawing with vertices placed on two parallel lines. The maximal $2$-layer 00-planar graphs (called caterpillars) are the maximal pathwidth-1 graphs [17, 18]. Angelini et al. [2] showed that $2$-layer $k$-planar graphs have pathwidth at most $k+1$ and gave a lower bound of $(k+3)/2$ for every odd number $k$. Hence, their result can be considered as a smooth generalization of the caterpillar result. Similarly, our treewidth bound $21.5k+$ 2 for outer $k$-planar graphs smoothly generalizes the treewidth-$2$ bound for maximal outerplanar graphs.
@@ 3 Triangulations
In this section, we present two simple strategies to triangulate maximal outer $k$-planar graphs. These triangulations will serve as tools to construct tree decompositions and balanced separators in Section 4. We assume that some outer $k$-planar drawing $\Gamma$ of a graph $G$ on $n$ vertices is given. During the triangulation procedures, we will modify the drawing, but we will never change the cyclic order of the vertices. Thus, let $v_{1},v_{2},\dots,v_{n}$ be the vertices of $G$ in the cyclic (counterclockwise) order defined by $\Gamma$. As we focus on maximal graphs, by Section 2, we have that $\langle v_{1},v_{2},\dots,v_{n},v_{1}\rangle$ is a cycle in $G$. Our goal is to construct a triangulation $\mathcal{T}$ of this cycle such that the edges of $\mathcal{T}$ cross only a limited number of edges of $G$. We refer to the edges of $\mathcal{T}$ as links in order to distinguish them from the edges of $G$. Every edge of the outer cycle is a link in $\mathcal{T}$; we call these links outer links. We will select a set of $n-3$ other pairwise non-intertwined pairs of vertices as inner links that, together with the outer links, form a triangulated $n$-gon. Note that some of the inner links may coincide with edges of $G$. We say that a link is pierced by an edge of $G$ if their endpoints are intertwined in the cyclic order. The piercing number of a link is the number of edges of $G$ that pierce the link. In particular the piercing number of outer links is 0, and if a link coincides with an edge, then its piercing number is at most $k$ by the outer $k$-planarity of $\Gamma$. We define the edge piercing number of $\mathcal{T}$ to be the maximum piercing number of any link of $\mathcal{T}$. In Section 4 we show how to use triangulations with small edge piercing number.
Our splitting procedure starts by declaring the link $\{v_{1},v_{n}\}$ active and considers all other vertices to lie on the right side of the link. See Fig. 1 for illustration. Clearly, the link $\{v_{1},v_{n}\}$ is not pierced. In each recursive step, the input is an active link $\{v_{i},v_{k}\}$ with $1\leq i<k\leq n$ (promised to be pierced at most some limited number of times), along with a distinguished right side $v_{i+1},\dots,v_{k-1}$, where currently no inner links have been selected apart from $\{v_{i},v_{k}\}$. (Vertices $v_{1},\dots,v_{i-1},v_{k+1},\dots,v_{n}$ form the left side.) The goal is to pick, among the vertices on the right side, a split vertex $v_{j}$ such that the new links $\{v_{i},v_{j}\}$ and $\{v_{j},v_{k}\}$ are pierced by at most some limited number of edges. The new links are added to $\mathcal{T}$ (completing the triangle $v_{i}v_{j}v_{k}$) and become active. The link $\{v_{i},v_{k}\}$ ceases to be active. The split gives rise to two new splitting instances $\{v_{i},v_{j}\}$ and $\{v_{j},v_{k}\}$, which are then solved recursively.
@@
We draw the vertices of $G$ on the unit circle, keeping the cyclic order defined by $\Gamma$. We set $\varepsilon>0$ to a sufficiently small value. Then we place $v_{i}$ at $(0,-1)$, $v_{k}$ at $(0,1)$, all the vertices from the left side of the link within arc distance $\varepsilon$ of $(-1,0)$ and spread the vertices from the right side of the link evenly along the right semicircle, see Fig. 2.
@@
Lemma 3.1 gives us an equivalent drawing of $G$ in which piercing edges cross the active link in the same order from bottom to top as their right endpoints in the cyclic order. This allows us to draw the piercing edges without crossings on the right side of the active link.
@@
We prove the lemma by providing an appropriate splitting procedure that recursively constructs the triangulation. Let $\Gamma$ be an outer $k$-planar drawing of $G$. We prove by induction on the number of steps of the splitting procedure that every active link has piercing number at most $12k-$ 1. The claim is true in the beginning, as we start with the link $\{v_{1},v_{n}\}$, which is not pierced at all.
@@
In the first case, $\lambda$ is not pierced by any edge of $G$. Observe that $\{v_{i},v_{i+1}\}$ is an edge, so $u$ has at least one neighbor in the right side of $\lambda$ in $G$. We select the split vertex, $w$, to be the neighbor of $u$ from the right side of $\lambda$ that is closest to $v$ in the circular order; see Fig. 3(a). In this case, the link $\{u,w\}$ coincides with an edge of $G$, and thus is pierced by at most $k$ edges. To see that the link $\{w,v\}$ is also pierced by at most $k$ edges, let $c^{\prime\prime}$ be a curve that starts in $w$, goes along $\{w,u\}$ towards $u$ on the right side (in the direction of walking), then follows $\{u,v\}$ on the right side, and ends at $v$. This curve crosses only the edges that pierce the link $\{u,w\}$, so at most $k$ edges pierce the link $\{w,v\}$ by Section 2.
In the second case, $\lambda$ is pierced by edges of $G$. Using Lemma 3.1, we modify the drawing $\Gamma$ so that the piercing edges cross $\lambda$ in the same bottom-to-top order as their right endpoints. Let $e_{1},e_{2},\dots,e_{\ell}$ denote the piercing edges and $w_{1},w_{2},\dots,w_{\ell}$ be their respective, not necessarily different, right endpoints in the order from bottom to top. By induction, we have $\ell\leq 2k-1$. Then we pick the middle edge $e_{j}$, where $j=\lceil(\ell+1)/2\rceil$, and set the split vertex $w=w_{j}$.
@@
Figure 3: Constructing a triangulation $\mathcal{T}$ with edge piercing number $12k-$ 1.
To bound the number of edges that pierce the link $\{u,w\}$, we draw the curve $c^{\prime}$ as in Fig. 3(b). We start the curve $c^{\prime}$ at $u$ and follow $\lambda$ until before it crosses $e_{j}$, then follow $e_{j}$ until we arrive at $w$. The curve $c^{\prime}$ has crossings with at most $\lceil(\ell-1)/2\rceil$ edges that pierce $\lambda$, and at most $k$ edges that have a crossing with $e_{j}$. Thus, there are at most $\lceil(\ell-1)/2\rceil+k\leq\lceil(2k-2)/2\rceil+k=2k-1$ edges that pierce the link $\{u,w\}$. We can argue symmetrically for the link $\{w,v\}$; see curve $c^{\prime\prime}$ in Fig. 3(b).
As we have bounded the piercing number of every active link during the triangulation by $12k-$ 1, we get the desired bound of $12k-$ 1 for the edge piercing number of the constructed triangulation.
Next, we present our refined triangulation method, where we consider an additional edge to get better bounds for the edge piercing number.
@@
The proof is similar to the proof of Lemma 3.3 and we use the same terminology. We present a splitting procedure and show that every active link is pierced by at most $k$ edges. For the base of induction, the link $\{v_{1},v_{n}\}$ has no piercing edges. Now, let $\lambda=\{v_{i},v_{k}\}$ be the active link and set $u=v_{i}$ and $v=v_{k}$. If $\lambda$ has no piercing edges, we proceed as in the proof of Lemma 3.3 and pick the split vertex to be the neighbor of $u$ from the right side of $\lambda$ that is closest to $v$. Otherwise, $\lambda$ is pierced by edges of $G$. Using Lemma 3.1, we assume that the piercing edges cross the active link in the same bottom-to-top order as their right endpoints. Now, let $e_{1},e_{2},\ldots,e_{\ell}$ be the piercing edges. We pick the middle edge $e_{j}$ with $j=\lceil(\ell+1)/2\rceil$. If $e_{j}$ has no more crossings on the right of the active link, then we pick the right endpoint of $e_{j}$ to be the split vertex. Both curves that start in the right endpoint of $e_{j}$, follow $e_{j}$ to $e$, and then follow $\lambda$ to either $u$ or $v$ have at most $\lceil(\ell-1)/2\rceil\leq k$ crossings each, so we can proceed. Now, let $x$ be the first crossing on the edge $e_{j}$ that occurs on the right side of $\lambda$, and let $\hat{e}$ denote the edge that crosses $e_{j}$ at $x$; see Fig. 4(a).
@@
We bound the piercing number of the link $\{v,w\}$. We start the curve $c^{\prime\prime}$ at $v$ and follow the link $\lambda$ until just before it crosses $e_{j}$, then follow $e_{j}$ until just before it crosses $\hat{e}$, then follow $\hat{e}$ until we arrive at $w$; see Fig. 4(a). In the first part, the curve crosses at most $\lceil(\ell-1)/2\rceil\leq\lceil(k-1)/2\rceil$ edges. The curve does not cross any edge in the middle part, as the crossing between $e_{j}$ and $\hat{e}$ is the first crossing along $e_{j}$ starting from $\lambda$. In the last part, the curve crosses at most $\min\{a,b\}\leq\lfloor(k-1)/2\rfloor$ edges. Thus, by Section 2, the piercing number of the link $\{v,w\}$ is at most $\lceil(k-1)/2\rceil+\lfloor(k-1)/2\rfloor\leq k-1$.
To bound the piercing number of the link $\{u,w\}$, we start the curve $c^{\prime}$ at $u$ and follow the link $\lambda$ until just after it crosses $e_{j}$ and then follow the curve $c^{\prime\prime}$ to $w$, see Fig. 4(a). This time, the number of crossings on the curve $c^{\prime}$ is at most $k$. Thus, the piercing number of the link $\{u,w\}$ is at most $k$.
Now we consider the last case, where $\hat{e}$ and $\lambda$ have a common vertex. In this case, we can choose the other vertex of $\hat{e}$ as $w$; see Fig. 4(b). Without loss of generality, we assume that $u$ is the common vertex. The edge $\{u,w\}$ has at most $k$ crossings by $k$-planarity. To bound the piercing number of the link $\{v,w\}$, observe that the edge $\{u,w\}$ has exactly $j$ crossings on the curve from $v$ to $x$ (including $x$). Therefore, as we did in the above, the edge $\{v,w\}$ is crossed by at most $(\ell-j)+(k-j)=k-(2j-\ell)$ edges. Note that $2j-\ell$ is $1$ if $\ell$ is $odd$ and 00 otherwise.
@@
Lastly, we combine the two strategies from Lemmas 3.3 and 3.5 to obtain a triangulation method for outer min-$k$-planar graphs.
@@
We proceed as in the proof of Lemma 3.3. We want to give a splitting procedure for which every active link has piercing number at most $12k-$ 1. Consult Fig. 3 for a reminder of the notation and the strategy. As $\Gamma$ is now an outer min-$k$-planar drawing, edges may have more than $k$ crossings. We call such edges heavy. Other edges, i.e., edges with no more than $k$ crossings, are light. By definition, heavy edges cross only light edges.
The first thing to notice is that as we allow for heavy edges now, the previous proof breaks in both cases when the curves $c^{\prime}$, $c^{\prime\prime}$ are routed along heavy edges. In particular, in the first case, $\lambda$ has no piercing edges, and we select the split vertex $w$ to be the neighbor of $u$ from the right side of $\lambda$ that is closest to $v$. If edge $\{u,w\}$ is heavy, then we cannot bound the number of crossings along the curve $c^{\prime\prime}$. In the second case, we select the split vertex to be the right endpoint $w_{j}$ of the middle piercing edge $e_{j}$. If edge $e_{j}$ is heavy, then we cannot bound the numbers of crossings along $c^{\prime}$ and $c^{\prime\prime}$.
To resolve the issue with the second case, we apply the strategy from the proof of Lemma 3.5. Observe that if $e_{j}$ is heavy, then for the first crossing $x$ with an edge $\hat{e}$ on the right side of $\lambda$, $\hat{e}$ is a light edge. Similarly, as in the proof of Lemma 3.5, see Fig. 4, we select the split vertex $w$ to be the endpoint of $\hat{e}$ that is different than $u$ and $v$ and in case both endpoints are, the one on the side of $x$ with fewer crossings. The piercing numbers of new links are at most $\lceil(2k-1)/2\rceil+k-1\leq 2k-1$.
To resolve the issue with the first case, we use a similar strategy. When edge $\{u,w\}$ is heavy, let $x$ to be the first crossing on $\{u,w\}$ from $u$ with an edge $\hat{e}$. Notice that $\hat{e}$ is light. We select the split vertex $w$ to be any right endpoint of $\hat{e}$. The curves that start at $w$, go along $\hat{e}$ to $x$, then along $\{u,w\}$ to $u$ and optionally to $v$ along $\lambda$ have at most $k$ crossings, only with edges that also cross $\hat{e}$.
@@
Each step of the splitting procedures used in Lemmas 3.33.11 can be implemented to run in $O(k)$ time, and the corresponding triangulations can be constructed in $O(nk)$ time.

### v2

@@ Abstract
treewidth, outerplanar graphs, outer $k$-planar graphs, outer min-$k$-planar graphs, cop number, separation number
@@ 1 Introduction
Treewidth is an important structural parameter because the running time of many graph algorithms and algorithms for drawing graphs depend on the treewidth. Recently treewidth-based techniques have been successfully applied for convex drawings, that is, straight-line drawings where vertices lie on a circle. For example, Bannister and Eppstein [8] ([7]) showed that if a graph admits a convex drawing with at most $k$ crossings in total, then its treewidth is bounded by a function of $k$. Furthermore, they used this fact to give, for a fixed $k$, a linear-time algorithm (via extended monadic second-order logic and Courcelle’s theorem [13]) to decide whether a convex drawing with at most $k$ crossings exists for a given graph. Chaplick et al. [12] ([11]) generalized this to bundled crossings (a crossing of two bundles of edges such that in each bundle the edges travel in parallel, which is counted as one crossing). Another prominent class of graphs with convex drawings is the class of outer $k$-planar graphs, that is, graphs that admit a convex drawing where each edge is crossed at most $k$ times. Unlike $k$-planar graphs (without the restriction on the placement of vertices), the treewidth of outer $k$-planar graphs can be bounded by a function of $k$ only (see discussion below). Similarly to Bannister and Eppstein [8], Chaplick et al. [10] used this to test in linear time, for any fixed $k$, whether a given graph is full outer $k$-planar, i.e., whether it admits an outer $k$-planar drawing where no crossing appears on the boundary of the outer face.
@@
For constant $k$, better treewidth bounds are known. A folklore result is that outerplanar graphs ($k=0$) have treewidth at most $2$. Auer et al. [5, 6] ([4]) showed that maximal outer $1$-planar graphs are chordal graphs and, therefore, have treewidth at most $3$, which is tight.
@@
A structurally similar type of result is known for $2$-layer $k$-planar graphs. A $2$-layer $k$-planar graph is a bipartite graph that admits a $k$-planar straight-line drawing with vertices placed on two parallel lines. The maximal $2$-layer 00-planar graphs (called caterpillars) are the maximal pathwidth-1 graphs [17, 18]. Angelini et al. [2] ([3]) showed that $2$-layer $k$-planar graphs have pathwidth at most $k+1$ and gave a lower bound of $(k+3)/2$ for every odd number $k$. Hence, their result can be considered as a smooth generalization of the caterpillar result. Similarly, our treewidth bound $21.5k+$ 2 for outer $k$-planar graphs smoothly generalizes the treewidth-$2$ bound for maximal outerplanar graphs.
@@ 3 Triangulations
In this section, we present two simple strategies to triangulate maximal outer $k$-planar graphs. These triangulations will serve as tools to construct tree decompositions and balanced separators in Section 4. We assume that some outer $k$-planar drawing $\Gamma$ of a graph $G$ on $n$ vertices is given. During the triangulation procedures, we will modify the drawing, but we will never change the cyclic order of the vertices. Thus, let $v_{1},v_{2},\dots,v_{n}$ be the vertices of $G$ in the cyclic (counterclockwise) order defined by $\Gamma$. As we focus on maximal graphs, by Section 2, we have that $\langle v_{1},v_{2},\dots,v_{n},v_{1}\rangle$ is a cycle in $G$. Our goal is to construct a triangulation $\mathcal{T}$ of this cycle such that the edges of $\mathcal{T}$ cross only a limited number of edges of $G$. We refer to the edges of $\mathcal{T}$ as links in order to distinguish them from the edges of $G$. Every edge of the outer cycle is a link in $\mathcal{T}$; we call these links outer links. We will select a set of $n-3$ other pairwise non-intertwined pairs of vertices as inner links that, together with the outer links, form a triangulated $n$-gon. Note that some of the inner links may coincide with edges of $G$. We say that a link is pierced by an edge of $G$ if their endpoints are intertwined in the cyclic order. The piercing number of a link is the number of edges of $G$ that pierce the link. In particular, the piercing number of outer links is 0, and if a link coincides with an edge, then its piercing number is at most $k$ by the outer $k$-planarity of $\Gamma$. We define the edge piercing number of $\mathcal{T}$ to be the maximum piercing number of any link of $\mathcal{T}$. In Section 4 we show how to use triangulations with small edge piercing number.
Our splitting procedure starts by declaring the link $\{v_{1},v_{n}\}$ active and considers all other vertices to lie on the right side of the link. See Figure 1 for illustration. Clearly, the link $\{v_{1},v_{n}\}$ is not pierced. In each recursive step, the input is an active link $\{v_{i},v_{k}\}$ with $1\leq i<k\leq n$ (promised to be pierced at most some limited number of times), along with a distinguished right side $v_{i+1},\dots,v_{k-1}$, where currently no inner links have been selected apart from $\{v_{i},v_{k}\}$. (Vertices $v_{1},\dots,v_{i-1},v_{k+1},\dots,v_{n}$ form the left side.) The goal is to pick, among the vertices on the right side, a split vertex $v_{j}$ such that the new links $\{v_{i},v_{j}\}$ and $\{v_{j},v_{k}\}$ are pierced by at most some limited number of edges. The new links are added to $\mathcal{T}$ (completing the triangle $v_{i}v_{j}v_{k}$) and become active. The link $\{v_{i},v_{k}\}$ ceases to be active. The split gives rise to two new splitting instances $\{v_{i},v_{j}\}$ and $\{v_{j},v_{k}\}$, which are then solved recursively.
@@
We draw the vertices of $G$ on the unit circle, keeping the cyclic order defined by $\Gamma$. We set $\varepsilon>0$ to a sufficiently small value. Then we place $v_{i}$ at $(0,-1)$, $v_{k}$ at $(0,1)$, all the vertices from the left side of the link within arc distance $\varepsilon$ of $(-1,0)$ and spread the vertices from the right side of the link evenly along the right semicircle, see Figure 2.
@@
Lemma 3.1 gives us an equivalent drawing of $G$ in which piercing edges cross the active link in the same order from bottom to top as their right endpoints occur in the cyclic order. This allows us to draw the piercing edges without crossings on the right side of the active link.
@@
We prove the lemma by providing an appropriate splitting procedure that recursively constructs the triangulation. Let $\Gamma$ be an outer $k$-planar drawing of the given maximal outer $k$-planar graph $G$. We prove by induction on the number of steps of the splitting procedure that every active link has piercing number at most $12k-$ 1. The claim is true in the beginning, as we start with the link $\{v_{1},v_{n}\}$, which is not pierced at all.
@@
In the first case, $\lambda$ is not pierced by any edge of $G$. Observe that $\{v_{i},v_{i+1}\}$ is an edge, so $u$ has at least one neighbor on the right side of $\lambda$ in $G$. We select the split vertex, $w$, to be the neighbor of $u$ from the right side of $\lambda$ that is closest to $v$ in the circular order; see Figure 3(a). In this case, the link $\{u,w\}$ coincides with an edge of $G$, and thus is pierced by at most $k$ edges. To see that the link $\{w,v\}$ is also pierced by at most $k$ edges, let $c^{\prime\prime}$ be a curve that starts in $w$, goes along $\{w,u\}$ towards $u$ on the right side (in the direction of walking), then follows $\{u,v\}$ on the right side, and ends at $v$. This curve crosses only the edges that pierce the link $\{u,w\}$, so at most $k$ edges pierce the link $\{w,v\}$ by Section 2.
In the second case, $\lambda$ is pierced by edges of $G$. Using Lemma 3.1, we modify the drawing $\Gamma$ so that the piercing edges cross $\lambda$ in the same bottom-to-top order as their right endpoints. Let $e_{1},e_{2},\dots,e_{\ell}$ denote the piercing edges, and let $w_{1},w_{2},\dots,w_{\ell}$ denote their respective, not necessarily different, right endpoints in the order from bottom to top. By induction, we have $\ell\leq 2k-1$. Then we pick the middle edge $e_{j}$, where $j=\lceil(\ell+1)/2\rceil$, and set the split vertex $w=w_{j}$.
@@
Figure 3: Constructing a triangulation $\mathcal{T}$ with edge piercing number at most $12k-$ 1.
To bound the number of edges that pierce the link $\{u,w\}$, we draw the curve $c^{\prime}$ as in Figure 3(b). We start the curve $c^{\prime}$ at $u$ and follow $\lambda$ until before it crosses $e_{j}$, then follow $e_{j}$ until we arrive at $w$. The curve $c^{\prime}$ crosses at most $\lceil(\ell-1)/2\rceil$ edges that pierce $\lambda$ and at most $k$ edges that pierce $e_{j}$. Thus, there are at most $\lceil(\ell-1)/2\rceil+k\leq\lceil(2k-2)/2\rceil+k=2k-1$ edges that pierce the link $\{u,w\}$. We can argue symmetrically for the link $\{w,v\}$; see curve $c^{\prime\prime}$ in Figure 3(b).
As we have bounded the piercing number of every active link during the triangulation procedure by $12k-$ 1, we get the desired bound of $12k-$ 1 for the edge piercing number of the constructed triangulation.
Next, we present our refined triangulation method, where we consider an additional edge to get a better bound for the edge piercing number.
@@
The proof is similar to the proof of Lemma 3.3, and we use the same terminology. Let again $\Gamma$ be an outer $k$-planar drawing of the given maximal outer $k$-planar graph $G$. We present a splitting procedure and show that every active link is pierced by at most $k$ edges. For the base of the induction, observe that the link $\{v_{1},v_{n}\}$ has no piercing edges. Now, let $\lambda=\{v_{i},v_{k}\}$ be the active link and set $u=v_{i}$ and $v=v_{k}$. If $\lambda$ has no piercing edges in $\Gamma$, we proceed as in the proof of Lemma 3.3 and pick the split vertex to be the neighbor of $u$ from the right side of $\lambda$ that is closest to $v$. Otherwise, $\lambda$ is pierced by edges of $G$. Using Lemma 3.1, we assume that the piercing edges cross the active link in the same bottom-to-top order as their right endpoints. Now, let $e_{1},e_{2},\ldots,e_{\ell}$ be the piercing edges, and let $w_{1},w_{2},\dots,w_{\ell}$ be their respective, not necessarily different, right endpoints in the order from bottom to top. We pick the middle edge $e_{j}$ with $j=\lceil(\ell+1)/2\rceil$. If $e_{j}$ has no crossings on the right side of the active link, then we pick $w_{j}$ to be the split vertex. The two curves that start in $w_{j}$, follow $e_{j}$ to $e$, and then follow $\lambda$ to either $u$ or $v$ have at most $\lceil(\ell-1)/2\rceil\leq k$ crossings each, so we can proceed. Now, let $x$ denote the first crossing on the edge $e_{j}$ that occurs on the right side of $\lambda$, and let $\hat{e}$ denote the edge that crosses $e_{j}$ at $x$; see Figure 4(a).
@@
We bound the piercing number of the link $\{v,w\}$. To this end, we start tracing a curve $c^{\prime\prime}$ at $v$ and follow the link $\lambda$ until just before it crosses $e_{j}$, then we follow $e_{j}$ until just before it crosses $\hat{e}$, and then we follow $\hat{e}$ until we arrive at $w$; see Figure 4(a). In the first part, the curve crosses at most $\lceil(\ell-1)/2\rceil\leq\lceil(k-1)/2\rceil$ edges. The curve does not cross any edge in the middle part, as the crossing between $e_{j}$ and $\hat{e}$ is the first crossing along $e_{j}$ starting from $\lambda$. In the last part, the curve crosses at most $\min\{a,b\}\leq\lfloor(k-1)/2\rfloor$ edges. Thus, by Section 2, the piercing number of the link $\{v,w\}$ is at most $\lceil(k-1)/2\rceil+\lfloor(k-1)/2\rfloor\leq k-1$.
To bound the piercing number of the link $\{u,w\}$, we start tracing a curve $c^{\prime}$ at $u$, follow the link $\lambda$ until just after it crosses $e_{j}$ and then follow the curve $c^{\prime\prime}$ to $w$, see Figure 4(a). The curve $c^{\prime}$ crosses the same set of edges as the curve $c^{\prime\prime}$ plus the edge $e_{j}$, so $c^{\prime\prime}$ has at most $k$ crossings. Thus, the piercing number of the link $\{u,w\}$ is at most $k$.
Now we consider the last case, where $\hat{e}$ and $\lambda$ have a common endpoint. In this case, we choose the other endpoint of $\hat{e}$ as $w$; see Figure 4(b). We first assume that $u$ is the common endpoint. The edge $\{u,w\}$ is an edge of $G$, so it has at most $k$ crossings. To bound the piercing number of the link $\{v,w\}$, we argue with the curve $c^{\prime\prime}$ defined in the previous case. Observe that the edge $\{u,w\}$ has exactly $j$ crossings on the part from $u$ to $x$ (including $x$). Therefore, we can argue as above and conclude that the edge $\{v,w\}$ is crossed by at most $(\ell-j)+(k-j)=k-(2j-\ell)$ edges. Note that $2j-\ell=2\lceil(\ell+1)/2\rceil-\ell$ is $1$ if $\ell$ is odd and $2$ otherwise. Similarly, if $v$ is the common endpoint, we obtain the bound $(j-1)+(k-(\ell-j+1))=k+(2j-\ell)-2$ for the piercing number of the link $\{u,w\}$.
@@
Lastly, we combine the two triangulation strategies from the proofs of Lemmas 3.3 and 3.5 to obtain a triangulation method for outer min-$k$-planar graphs.
@@
We proceed as in the proof of Lemma 3.3. We describe a splitting procedure such that every active link has piercing number at most $12k-$ 1. Consult Figure 3 for a reminder of the notation and the strategy. As the given drawing is outer min-$k$-planar, edges may now have more than $k$ crossings. We call such edges heavy and the other edges light. By definition, heavy edges cross only light edges.
We first explain why the previous proof breaks in both cases if the curves $c^{\prime}$ and $c^{\prime\prime}$ are routed along heavy edges. Recall that in the first case when $\lambda$ has no piercing edges, we selected the split vertex $w$ to be the neighbor of $u$ from the right side of $\lambda$ that is closest to $v$. If, however, the edge $\{u,w\}$ is heavy, then we cannot bound the number of crossings along the curve $c^{\prime\prime}$. In the second case, we selected the split vertex to be the right endpoint $w_{j}$ of the middle piercing edge $e_{j}$. If the edge $e_{j}$ is heavy, then we cannot bound the numbers of crossings along $c^{\prime}$ and $c^{\prime\prime}$.
To resolve the issue with the second case, we apply the strategy from the proof of Lemma 3.5. Observe that if $e_{j}$ is heavy, then for the first crossing $x$ with an edge $\hat{e}$ on the right side of $\lambda$, the edge $\hat{e}$ is a light. Similarly, as in the proof of Lemma 3.5 (see Figure 4), we select the split vertex $w$ to be the endpoint of $\hat{e}$ that is different from $u$ and $v$ and in case both endpoints are, the one on the side of $x$ with fewer crossings. The piercing numbers of the two new links are at most $\lceil(2k-1)/2\rceil+k-1\leq 2k-1$.
To resolve the issue with the first case, we use a similar strategy. If the edge $\{u,w\}$ is heavy, let $x$ be the crossing on $\{u,w\}$ that is closest to $u$ and let $\hat{e}$ be the crossing edge. Note that $\hat{e}$ is light. We select the split vertex $w$ to be any right endpoint of $\hat{e}$. Now consider the two curves that start at $w$, go along $\hat{e}$ to $x$, then along $\{u,w\}$ to $u$, and optionally to $v$ along $\lambda$. Both curves have at most $k$ crossings, only with edges that also cross $\hat{e}$.
@@
Given the intersection graph of the edges of $G$, the triangulations in Lemmas 3.33.11 can be constructed in $O(nk)$ time. Each step of the splitting procedures presented in Lemmas 3.33.11 can be implemented to run in $O(k)$ time, as the edges that pierce the new active links also pierce at least one of the edges $\lambda$, $e_{j}$, and $\hat{e}$.